# Optical computing using telecom device platform for nonlinear mapping to high-dimensional feature space

### v1

@@ 3 Benchmarks
For the classification task, we used the standard MNIST dataset, splitting it into training and testing subsets with an 80/20 ratio. Since linear classifiers already achieve around 90% accuracy on the original 784-pixel images, we increased the task difficulty by downsampling the images to just 15 pixels. We accomplished this using column-pivoting QR decomposition to identify and retain the most informative pixels. To ensure a proper evaluation, this decomposition was performed exclusively on the training subset to prevent any information from the testing subset from influencing the pre-processing stage. The identified pixel positions were subsequently used to downsample images in the testing subset. Applying this approach, we observed an accuracy of 42% on the testing subset using linear classification. The 15 selected pixels were then input into the ELM to achieve a high-dimensional nonlinear mapping. A linear classifier was trained on the mapped samples from the training subset within this higher-dimensional feature space. The performance of the classifier was then evaluated using the mapped samples from the testing subset. When using WDM for encoding, we simultaneously transmit different pixels across separate spectral channels.
@@ 4 Results and discussion
We used symbol-to-pulse train encoding with Gaussian, skewed Gaussian, or a trainable mask. The input pulse had a peak power of 50 W. Once more, we would like to stress that this power corresponds to NOLm based on SMF and can be substantially scaled down by using highly nonlinear fiber, or integrated MZI/NOLM with other material platform. When employing the trainable mask, it was first set digitally with ideal (infinitely steep) transitions, then simulated for encoding by an electro-optical modulator with a 20 GHz analog bandwidth. For WDM multi-channel spectral encoding, we used a frequency step of $\Delta\nu=50~{}\text{GHz}$ between channels, with channels number 1, 2, 3, … corresponding to frequency shifts of 0, $-\Delta\nu$, $+\Delta\nu$, $-2\Delta\nu$, … For time series forecasting, we encoded time-shifted copies of the same symbol sequences in other spectral channels so that channel 2 contains a 1-symbol delayed copy of the channel 1 sequence, channel 3 contains a 2-symbol delayed copy of the channel 1 sequence, etc. Using WDM for the classification task, we simply encoded different pixels into different WDM channels in parallel, thus reducing the number of sequential symbols to be fed into the ELM.
The NOLM was characterized using parameters of a single mode fiber: a nonlinear coefficient $\gamma=10.8\times 10^{-3}~{}\mathrm{W^{-1}\,m^{-1}}$, group velocity dispersion coefficient $\beta_{2}=26\times 10^{-3}~{}\mathrm{ps^{2}\,m^{-1}}$, fiber length $L_{\mathrm{NOLM}}=100~{}\text{m}$, and the coupling ratio $\kappa=0.3$. Evidently, employing highly nonlinear fibers or other material platforms, all resulting optical pulse parameters can be easily scaled.
For the time series forecasting task, the model was trained using 4,000 symbols and tested on 100-1000 symbols depending on the chaoticity of the time series. We used batches of 15 symbols to predict the next one. To make a multi-step prediction, we utilized the autoregressive approach, when the newly predicted symbols are used to predict the next ones. A regularization amplitude was optimized for performance and to prevent overfitting. For the classification task, the model was trained using 5,000 images from the training set, which were sub-sampled using column-pivoting QR decomposition. The pixel locations chosen for the training subset were then applied to the testing subset during the testing phase, which also included 5,000 sample images.
@@
Figure 6: Performance comparison of the linear model on time-series prediction tasks. Poor prediction accuracy is achieved for all the Mackey-Glass (a) and Rössler (b) time series, which indicates the inability of the linear model to capture the complex dynamics of these systems.
Although the single-step prediction normalized mean squared error (NMSE) appeared relatively low, at $\text{NMSE}_{\text{MG}}\approx 5\cdot 10^{-6}$, $\text{NMSE}_{\text{R}}\approx 2\cdot 10^{-4}$, the model failed at long-term prediction. This failure was due to the linear system’s inability to capture and replicate the underlying dynamics of these complex systems.
@@
Figure 8: Example of autoregressive prediction of Mackey-Glass series for 1000 symbols using PELM
The normalized mean squared error (NMSE) averaged over multiple predicted sequences, similar to shown in Fig. 8 but taken from different parts of the Mackey-Glass sequence, is equal to $\text{NMSE}=0.097$ for the autoregressive prediction depth of 1000 symbols, which is a significant improvement compared to the purely linear model, shown in Fig. 6(a).
@@
Figure 10: Normalized mean square error (NMSE) as a function of autoregressive prediction depth for different encoding techniques: (a) single-channel Gaussian pulse encoding, (b) 5-channel WDM with Gaussian pulse encoding, and (c) single-channel trained encoding mask. Errors for individual predicted sequences are shown in gray, and the average error across all predicted sequences is highlighted in red.
@@
Figure 12: Normalized Mean Squared Error (NMSE) versus autoregressive prediction depth for the Rössler time series (a); the results demonstrate reduced prediction error when using 3 WDM channels (b) or a trainable encoding mask (c), compared to single-channel encoding.
@@
We also tested the proposed approach on SOA as a nonlinear mapping device. All system parameters are kept the same except the peak power $P_{0}=0.7~{}\text{W}$. SOA parameters used for simulation (see eqs. 5-6), without loss of generality, are the following: $\beta=5$, small signal gain $30$ dB, $\tau_{c}=200$ ps, $E_{sat}=8$ pJ. Figure 13 shows how applying WDM encoding (b) and trainable encoding mask (c) outperform the standard Gaussian encoding mask (a) for Rössler attractor prediction task.
Figure 13: Normalized Mean Squared Error (NMSE) versus autoregressive prediction depth when using SOA as a nonlinear mapping device, for the Rössler time series (a); the results demonstrate reduced prediction error when using 3 WDM channels (b) or a trainable encoding mask (c), compared to single-channel encoding.

### v2

@@ 3 Benchmarks
To solve Eq. 14, we used the dde23 MATLAB solver with an adaptive step size to ensure target relative and absolute tolerances of RelTol=1e-6 and AbsTol=1e-8. The solution was then interpolated onto a uniform grid with a step size of 1, following the approach in Jaeger and Haas [60]. For the Rössler attractor, we used the ode45 MATLAB solver with adaptive step to ensure the same relative and absolute tolerances. We then interpolated the solution onto a uniform grid with a step size of 0.25. For the Rössler attractor, the time series was rescaled to lie within the range $[0.1,1.1]$ to avoid zero or negative values, which are incompatible with optical power levels in our simulation. In contrast, the Mackey–Glass time series was used without normalization, as its amplitude remained within a suitable range for our modeling.
@@
To validate our Lyapunov time estimation based on trajectory divergence, we also computed the Lyapunov exponent for the Rössler attractor using the standard variational method based on linearization of the system equations. Specifically, we integrated both the original system and its linearized form (variational equations) and periodically orthonormalized the perturbation vectors using QR decomposition to estimate the largest Lyapunov exponent. This yielded a value for inverse Lyapunov exponent of approximately 13.4, which is consistent with the rate inferred from the divergence of nearby trajectories shown in Fig. 5. For the Mackey–Glass system, a similar analysis could not be performed due to its infinite-dimensional nature as a delay differential equation, which complicates the formulation and integration of variational equations. Therefore, for Mackey–Glass, we relied solely on the trajectory-based method to estimate the Lyapunov time.
@@
For the classification task, we took 10,000 samples from the MNIST dataset and split them 50/50 for training/testing. Since linear classifiers already achieve around 90% accuracy on the original 784-pixel images, we increased the task difficulty by downsampling the images to just 15 pixels. We accomplished this using column-pivoting QR decomposition to identify and retain the most informative pixels. To do this, we reshaped the MNIST dataset into a matrix in the form 10,000 by 784 and applied column-pivoting QR decomposition, following the data-driven QR sensing paradigm, which is basically a compressed sensing approach performed on a tailored basis [61]. The pivoting QR factorization seeks to reorder the columns such that the most ”informative” columns are moved to the forefront. The measure of ”informativeness” here is rooted in the magnitudes of the diagonal entries of the upper triangular matrix $\mathbf{R}$ resulting from the QR factorization. Columns, or, in our case, pixels, corresponding to larger diagonal entries in $\mathbf{R}$ are deemed more significant. Thus, we select the most informative columns. To ensure a proper evaluation, this decomposition was performed exclusively on the training subset to prevent any information from the testing subset from influencing the pre-processing stage. The identified pixel positions were subsequently used to downsample images in the testing subset. We used Linear Discriminant Analysis (LDA) for multi-class classification using either original selected pixels (linear classification) or processed by the photonic ELM. Applying this approach, we observed an accuracy of 42% on the testing subset using linear classification. The 15 selected pixels were then input into the ELM to achieve a high-dimensional nonlinear mapping. A linear classifier was trained on the mapped samples from the training subset within this higher-dimensional feature space. The performance of the classifier was then evaluated using the mapped samples from the testing subset. When using WDM for encoding, we simultaneously transmit 5 different pixels across 5 separate spectral channels, effectively sending the input in batches of 5 pixels at a time, with 3 such batches per image.
@@ 4 Results and discussion
We used symbol-to-pulse train encoding with Gaussian, skewed Gaussian, or a trainable mask. The input pulse had a peak power of 50 W. Once more, we would like to stress that this power corresponds to NOLM based on SMF and can be substantially scaled down by using highly nonlinear fiber, or integrated MZI/NOLM with other material platform. When employing the trainable mask, it was first set digitally with ideal (infinitely steep) transitions, then simulated for encoding by an electro-optical modulator with a 20 GHz analog bandwidth. For WDM multi-channel spectral encoding, we used a frequency step of $\Delta\nu=50~{}\text{GHz}$ between channels, with 5 channels corresponding to frequency shifts of 0, $-\Delta\nu$, $+\Delta\nu$, $-2\Delta\nu$, and $+2\Delta\nu$. For time series forecasting, we encoded time-shifted copies of the same symbol sequences in other spectral channels so that channel 2 contains a 1-symbol delayed copy of the channel 1 sequence, channel 3 contains a 2-symbol delayed copy of the channel 1 sequence, etc. A similar approach was used for memory enhancement in RC in microresonators [37]. However, unlike the approach in [37], our scheme additionally employs trainable encoding. Theoretical studies [62] have also explored memory enhancement via input-stage modifications, particularly in SOAs and microring resonators. Our work extends these insights by demonstrating an implementation in a WDM-based framework, offering flexible spectral encoding, trainable masking for dynamic input modulation, and a scalable approach using commercially available telecom components. We did not demultiplex the WDM channels at the output and instead simulated the readout as being measured directly from the nonlinear transform’s output using a single photodiode. Using WDM for the classification task, we simply encoded different pixels into different WDM channels in parallel, thus reducing the number of sequential symbols to be fed into the ELM.
The NOLM was characterized using parameters of a single mode fiber: a nonlinear coefficient $\gamma=0.8\times 10^{-3}~{}\mathrm{W^{-1}\,m^{-1}}$, group velocity dispersion coefficient $\beta_{2}=26\times 10^{-3}~{}\mathrm{ps^{2}\,m^{-1}}$, fiber length $L_{\mathrm{NOLM}}=100~{}\text{m}$, and the coupling ratio $\kappa=0.3$. Evidently, employing highly nonlinear fibers or other material platforms, all resulting optical pulse parameters can be easily scaled.
For the time series forecasting task, the model was trained using 4,000 symbols and tested on 100-1000 symbols depending on the chaoticity of the time series. We used batches of $N=15$ symbols to predict the next one. To make a multi-step prediction, we utilized the autoregressive approach, when the newly predicted symbols are used to predict the next ones. A regularization amplitude was optimized for performance and to prevent overfitting. For the classification task, the model was trained using 5,000 images from the training set, which were sub-sampled using column-pivoting QR decomposition. The pixel locations chosen for the training subset were then applied to the testing subset during the testing phase, which also included 5,000 sample images.
@@
Figure 6: Performance, for comparison, of the linear model on time-series prediction tasks. Poor prediction accuracy is achieved for all the Mackey-Glass (a) and Rössler (b) time series, which indicates the inability of the linear model to capture the complex dynamics of these systems.
Although the single-step prediction NMSE appeared relatively low, at $\text{NMSE}_{\text{MG}}\approx 5\cdot 10^{-6}$, $\text{NMSE}_{\text{R}}\approx 2\cdot 10^{-4}$, the model failed at long-term prediction. This failure was due to the linear system’s inability to capture and replicate the underlying dynamics of these complex systems.
@@
Figure 8: Example of autoregressive prediction of Mackey-Glass series for 1000 symbols using PELM.
NMSE averaged over multiple predicted sequences, similar to shown in Fig. 8 but taken from different parts of the Mackey-Glass sequence, is equal to $\text{NMSE}=0.097$ for the autoregressive prediction depth of 1000 symbols, which is a significant improvement compared to the purely linear model, shown in Fig. 6(a).
@@
Compared to other photonic RC approaches for Mackey–Glass prediction, our method outperforms [63], achieving an NMSE of 0.01 for predicting 300 symbols (six quasi-periods), whereas [63] reports an NMSE of 0.1. In comparison to the delay-based RC in [26], our approach achieves similar accuracy for single-step prediction; however, a direct comparison of multi-step prediction performance is not possible due to the lack of reported error metrics in [26].
Figure 10: NMSE as a function of autoregressive prediction depth for different encoding techniques: (a) single-channel Gaussian pulse encoding, (b) 5-channel WDM with Gaussian pulse encoding, and (c) single-channel trained encoding mask. Errors for individual predicted sequences are shown in gray, and the average error across all predicted sequences is highlighted in red.
@@
Figure 12: NMSE versus autoregressive prediction depth for the Rössler time series (a); the results demonstrate reduced prediction error when using 3 WDM channels (b) or a trainable encoding mask (c), compared to single-channel encoding.
@@
We also tested the proposed approach on SOA as a nonlinear mapping device. All system parameters are kept the same except the peak power $P_{0}=0.7~{}\text{W}$. SOA parameters used for simulation (see eq. 5) are the following: $\beta=5$, small signal gain $30$ dB, $\tau_{c}=200$ ps, $E_{sat}=8$ pJ. All time series and computing parameters are identical to those in Fig. 12, except that the nonlinear transform is now performed using an SOA. Figure 13 shows how applying WDM encoding (b) and trainable encoding mask (c) outperform the standard Gaussian encoding mask (a) for Rössler attractor prediction task.
Figure 13: NMSE versus autoregressive prediction depth when using SOA as a nonlinear mapping device, for the Rössler time series (a); the results demonstrate reduced prediction error when using 3 WDM channels (b) or a trainable encoding mask (c), compared to single-channel encoding.